OpenStudy (anonymous):

solve the proportion using cross products 18/20 = k/110 A.99 B.122 C.3.3 D.2.9

OpenStudy (anonymous):

\[\frac{ 18 }{ 20}=\frac{ k }{ 110 }\] \[20k = 18 (110)\] So the answer is A.

OpenStudy (anonymous):

thank you